Find the eccentricity, foci, and directrices of the ellipse 2x² + y2 = 4.
Answer #1

solution Given that 2x+y=4 2x² tye al 4 Xla ge + comparto de 2 + 2 = 1 acb. ar eccentricity e- Inal 62 5²=4 e = 1-2 4 Shr Foc